I have a question that is giving me trouble, if anyone could help me out it would be appreciated.
Nick has two cans of white paint, one with a 2% yellow pigment and the other with a 5% yellow pigment. Nick wants to mix the two paints to get paint with a 4% yellow pigment. How much of the 5% yellow pigment paint should be mixed with 0.4 gallons of the 2% yellow pigment paint to get the desired paint?

Answer by stanbon(75887)

Nick has two cans of white paint, one with a 2% yellow pigment and the other with a 5% yellow pigment.
Nick wants to mix the two paints to get paint with a 4% yellow pigment.
How much of the 5% yellow pigment paint should be mixed with 0.4 gallons of the 2% yellow pigment paint to get the desired paint?
------------------------
Amt. of 5% paint = x ; Amt. of yellow = 0.05x gallons
----------------------------
Amt. of 2% paint = 0.4 gals : Amt. of yellow = 0.02*0.4 = 0.008 gals
----------------------------------
EQUATION:
yellow + yellow = yellow
0.05x + 0.008 = 0.04(x+.4)
50x + 8 = 40x + 16
10x = 8
x = 0.8 gallons (Amount of 5% point that must be added)
====================================================
Cheers,
Stan H.